# Spectral Differential Network Analysis for High-Dimensional Time Series

### v1

@@ A.4 Simulation Specifics and Results
To solve the SDD estimation problem (1), the alternating direction method of multipliers (ADMM) algorithm from Yuan et al. (2017) was used. Specifically, we used the L1_dts function in the Difdtl package available on GitHub at SusanYuan/Difdtl (GPL ($\geq$ 2)). To generate the sequence of penalties, $\{\tau_{n_{1},n_{2}}\}$, we used 20 values on a log-linear scale from $0.001*\tau_{n_{1},n_{2},\mathrm{max}}$ to $\tau_{n_{1},n_{2},\mathrm{max}}$ where $\tau_{n_{1},n_{2},\mathrm{max}}$ represents the minimum value where all entries of $\hat{\Delta}$ are 0. In this case, $\tau_{n_{1},n_{2},\mathrm{max}}=2\max(|\hat{\Sigma}_{1}-\hat{\Sigma}_{2}|)$. The penalty $\tau_{n_{1},n_{2}}^{*}$ was chosen as the $\tau_{n_{1},n_{2}}$ that minimizes the BIC which is computed as
$\mathrm{BIC}(\hat{\Delta})=2n\left\|\frac{1}{2}\left(\hat{\Sigma}_{1}\hat{\Delta}\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\hat{\Delta}\hat{\Sigma}_{1}-\hat{\Sigma}_{2}+\hat{\Sigma}_{1}\right)\right\|_{F}+\log(2n)\|\hat{\Delta}\|_{0}\,.$
@@
1.01 (1.75) 0.65 (0.47) 0.04 (0.06) 1.00 (0.00) 1.01 (0.04)
0.96 (1.02) 0.66 (0.47) 0.03 (0.04) 1.00 (0.00) 0.96 (0.06)
7.33 (9.40) 0.83 (0.37) 0.27 (0.34) 1.00 (0.00) 0.69 (0.34)
@@
100 -597 (265) 0.39 (0.45) -0.46 (0.25) 0.05 (0.02) 0.01 (0.06)
200 -943 (431) 0.55 (0.47) -0.42 (0.20) 0.08 (0.04) -0.03 (0.07)
500 -1469 (606) 0.38 (0.44) -0.23 (0.21) 0.13 (0.05) -0.12 (0.18)
1000 -1842 (808) 0.39 (0.44) -0.21 (0.26) 0.16 (0.07) -0.20 (0.21)
2000 -2511 (1012) 0.35 (0.42) -0.14 (0.22) 0.21 (0.09) -0.16 (0.17)
Table 5: Simulation 1. Difference Between SDD and Hard Thresholding Method for Each Metric.
@@
200 133 (183) 0.19 (0.79) 0.35 (0.18) -0.01 (0.02) -0.72 (0.66)
500 304 (328) -0.26 (0.64) 0.61 (0.26) -0.02 (0.03) -0.33 (0.20)
1000 439 (514) -0.36 (0.61) 0.47 (0.37) -0.04 (0.04) -0.29 (0.28)
2000 561 (852) -0.46 (0.53) 0.45 (0.40) -0.05 (0.07) -0.22 (0.44)
@@
100 -255 (73.4) -0.02 (0.01) -0.42 (0.06) 0.02 (0.01) -0.02 (0.04)
200 -607 (457) -0.02 (0.02) -0.43 (0.05) 0.05 (0.04) -0.01 (0.07)
500 -3548 (562) 0.01 (0.05) -0.52 (0.15) 0.30 (0.05) -0.21 (0.27)
1000 -6463 (585) 0.28 (0.43) -0.66 (0.23) 0.55 (0.05) -0.61 (0.44)
2000 -9032 (1021) 0.08 (0.14) -0.58 (0.14) 0.77 (0.09) -1.57 (0.37)
Table 7: Simulation 2. Difference Between SDD and Hard Thresholding Method for Each Metric.
@@
200 1.90 (6.20) -0.06 (0.24) -0.00 (0.02) -0.00 (0.00) -2.05 (1.79)
500 17.3 (15.9) -0.45 (0.49) -0.02 (0.08) -0.00 (0.00) -0.10 (0.20)
1000 22.5 (39.3) -0.49 (0.58) 0.06 (0.13) -0.00 (0.00) 0.05 (0.11)
2000 112 (73.7) -0.77 (0.38) 0.28 (0.14) -0.01 (0.01) -0.16 (0.15)
@@
200 1.72 (1.50) 0.54 (0.50) 0.07 (0.05) 0.00 (0.00) -0.77 (0.91)
500 19.6 (29.7) -0.10 (0.59) 0.16 (0.18) -0.00 (0.00) -0.14 (0.14)
1000 12.9 (20.1) 0.13 (0.60) 0.25 (0.22) 0.00 (0.00) -0.17 (0.19)
2000 54.5 (58.5) -0.26 (0.44) 0.51 (0.37) -0.00 (0.00) -0.17 (0.33)
A.5 Additional $\mu\mathrm{ECoG}$ Analyses
Figure 12: Relationship Between Probability of Edge Change and Distance Between Electrodes Among Control Sessions. Lines are generated by smoothing using a GAM. Shaded areas represent 95% confidence intervals. Electrodes that are close together have a higher probability of edge change.
Figure 13: Relationship Between Average Distance of Electrode Pair to Closest Stimulation Site and Distance Between Electrodes Among All Sessions. Lines are generated by smoothing using a LOESS smoother. Shaded areas represent 95% confidence intervals. Electrodes pairs that are either very close or very far from stimulation sites are close together.
Figure 13 shows that electrode pairs that have a small or large mean stim distance also have small distance between them. This makes sense as if two electrodes are close together both are either far or close to the stim sites. Taken with Figure 12, this means that even in control sessions we expect a U-shaped relationship between distance to stimulation site and probability of edge change.

### v2

@@ Appendix D SIMULATION SPECIFICS AND RESULTS
To solve the SDD estimation problem (1), the alternating direction method of multipliers (ADMM) algorithm from Yuan et al. (2017) was used. Specifically, we used the L1_dts function in the Difdtl package available on GitHub at SusanYuan/Difdtl (GPL ($\geq$ 2)). To generate the sequence of penalties, $\{\tau_{n_{1},n_{2}}\}$, we used 20 values on a log-linear scale from $0.001*\tau_{n_{1},n_{2},\mathrm{max}}$ to $\tau_{n_{1},n_{2},\mathrm{max}}$ where $\tau_{n_{1},n_{2},\mathrm{max}}$ represents the minimum value where all entries of $\hat{\Delta}$ are 0. In this case, $\tau_{n_{1},n_{2},\mathrm{max}}=2\max(|\hat{\Sigma}_{1}-\hat{\Sigma}_{2}|)$. The penalty $\tau_{n_{1},n_{2}}^{*}$ was chosen as the $\tau_{n_{1},n_{2}}$ that minimizes the eBIC which is computed as
$\mathrm{eBIC}(\hat{\Delta})_{\gamma}=\min(n_{1},n_{2})\left\|\frac{1}{2}\left(\hat{\Sigma}_{1}\hat{\Delta}\hat{\Sigma}_{2}+\hat{\Sigma}_{2}\hat{\Delta}\hat{\Sigma}_{1}-\hat{\Sigma}_{2}+\hat{\Sigma}_{1}\right)\right\|_{\infty}+\log(\min(n_{1},n_{2}))|E|+4\gamma|E|\log(p)\,,$ (7)
where $|E|$ is the number of unique edges in $\hat{\Delta}$. Since $\hat{\Delta}$ represents the difference in expanded spectral densities $|E|$ is the number of non-zero entries in the upper triangular portions, including diagonals, of the submatrices $\hat{\Delta}_{1:p,1:p}$ and $\hat{\Delta}_{1:p,(p+1):2p}$. For all applications we use $\gamma=0.5$. For FGL, Naïve, and Hard Thresholding, 20 tuning parameter values were also used. FGL requires two tuning parameter values, $\lambda_{1},\lambda_{2},$ so a grid of 20 $(\lambda_{1},\lambda_{2})$ pairs was considered. Since the parameter $\lambda_{1}$ controls the sparsity of each inverse matrix while $\lambda_{2}$ controls sparsity of the difference we considered only two values for $\lambda_{1}$ and 10 values for $\lambda_{2}$. The values for $\lambda_{1}$ were $(0.01\lambda_{1,\max},0.1\lambda_{1,\max})$ where $\lambda_{1,\max}=\left\|\hat{\Sigma}_{1,O}+\hat{\Sigma}_{2,O}\right\|_{\infty}/2$ and $\hat{\Sigma}_{i,O}$ is the off-diagonals of $\hat{\Sigma}_{i}$. The $\lambda_{2}$ sequence was generated using 10 values on a log-linear scale from $0.0001*\lambda_{2,\max}$ to $\lambda_{2,\max}$ where $\lambda_{2,\max}=\max\left(\left\|\hat{\Sigma}_{1,O}\right\|_{\infty}-0.01\lambda_{1,\max},\left\|\hat{\Sigma}_{2,O}\right\|_{\infty}-0.01\lambda_{1,\max}\right)$. For the Naïve method, we first computed $\lambda_{\max,i}=\left\|\hat{\Sigma}_{i}\right\|_{\infty}$ in each condition and then computed the tuning parameters as 20 evenly-spaced values on a log-linear scale from $0.0001\lambda_{\max,i}$ to $\lambda_{\max,i}$. Similarly, for the Hard Thresholding method, we used 20 evenly-spaced values on a log-linear scale from $\min(|\hat{\Delta}_{H}|)$ to $\max(|\hat{\Delta}_{H}|)$ where $\hat{\Delta}_{H}$ is generated by inverting $\hat{\Sigma}_{i}$ and forming the difference. Note that for $n=100$, it is not possible to invert $\hat{\Sigma}_{i}$.
@@
0.72 (0.06) 0.65 (0.05) 0.03 (0.00) 1.00 (0.00) 1.01 (0.00)
0.81 (0.07) 0.66 (0.05) 0.03 (0.00) 1.00 (0.00) 0.97 (0.00)
2.67 (1.78) 0.65 (0.05) 0.04 (0.01) 1.00 (0.00) 0.96 (0.00)
Table 4: Simulation 1. Difference Between SDD and FGL for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 -2062 (194) 0.44 (0.06) -0.58 (0.03) 0.17 (0.02) 0.06 (0.01)
200 -2428 (137) 0.50 (0.04) -0.58 (0.02) 0.21 (0.01) 0.06 (0.02)
500 -4952 (326) 0.64 (0.03) -0.45 (0.02) 0.42 (0.03) 0.10 (0.02)
1000 -5931 (325) 0.63 (0.03) -0.33 (0.02) 0.51 (0.03) -0.03 (0.03)
2000 -6259 (329) 0.55 (0.03) -0.27 (0.02) 0.54 (0.03) -0.08 (0.03)
@@
100 -649 (46.0) 0.43 (0.06) -0.40 (0.03) 0.05 (0.00) 0.00 (0.00)
200 -1060 (56.4) 0.50 (0.04) -0.50 (0.01) 0.09 (0.00) -0.01 (0.01)
500 -1754 (71.9) 0.64 (0.03) -0.40 (0.02) 0.15 (0.01) -0.02 (0.01)
1000 -2258 (103) 0.62 (0.03) -0.33 (0.02) 0.19 (0.01) -0.07 (0.01)
2000 -3109 (149) 0.55 (0.03) -0.27 (0.02) 0.27 (0.01) -0.04 (0.01)
Table 6: Simulation 1. Difference Between SDD and Hard Thresholding for Each Metric.
@@
200 16.0 (1.00) 0.13 (0.07) 0.28 (0.02) -0.00 (0.00) -0.67 (0.06)
500 20.3 (1.15) -0.03 (0.06) 0.48 (0.02) 0.00 (0.00) -0.22 (0.01)
1000 26.1 (1.86) -0.06 (0.05) 0.51 (0.04) -0.00 (0.00) -0.27 (0.03)
2000 -182 (52.5) -0.05 (0.05) 0.37 (0.05) 0.02 (0.00) -0.50 (0.07)
Table 7: Simulation 2. Difference Between SDD and FGL for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 -4674 (227) -0.00 (0.00) -0.50 (0.02) 0.40 (0.02) -1.25 (0.16)
200 -5296 (185) -0.00 (0.00) -0.60 (0.02) 0.45 (0.02) -1.15 (0.12)
500 -10199 (119) -0.00 (0.00) -0.97 (0.01) 0.87 (0.01) -3.20 (0.07)
1000 -9727 (155) 0.29 (0.04) -0.84 (0.02) 0.83 (0.01) -2.10 (0.06)
2000 -9189 (177) 0.33 (0.04) -0.82 (0.02) 0.79 (0.02) -1.35 (0.05)
@@
100 -257 (10.7) -0.02 (0.00) -0.42 (0.01) 0.02 (0.00) -0.02 (0.01)
200 -608 (46.2) -0.02 (0.00) -0.43 (0.00) 0.05 (0.00) -0.01 (0.01)
500 -3564 (56.3) -0.00 (0.00) -0.54 (0.01) 0.30 (0.00) -0.20 (0.03)
1000 -6483 (58.4) 0.29 (0.04) -0.70 (0.02) 0.55 (0.00) -0.59 (0.04)
2000 -9139 (98.6) 0.33 (0.04) -0.82 (0.02) 0.78 (0.01) -1.34 (0.04)
Table 9: Simulation 2. Difference Between SDD and Hard Thresholding for Each Metric.
@@
200 0.18 (0.14) -0.06 (0.02) -0.00 (0.00) -0.00 (0.00) -2.05 (0.18)
500 1.31 (0.36) -0.46 (0.05) -0.04 (0.01) -0.00 (0.00) -0.09 (0.02)
1000 2.87 (0.48) -0.49 (0.06) 0.02 (0.01) -0.00 (0.00) 0.05 (0.01)
2000 5.44 (0.84) -0.51 (0.06) 0.06 (0.01) -0.00 (0.00) 0.05 (0.01)
Table 10: Simulation 3. Difference Between SDD and FGL for Each Metric.
n # Est edges Precision Recall Accuracy RRMSE
100 -2948 (157) 0.81 (0.05) -0.76 (0.03) 0.25 (0.01) -0.01 (0.01)
200 -2979 (96.9) 0.72 (0.04) -0.75 (0.02) 0.25 (0.01) -0.01 (0.02)
500 -7711 (254) 0.63 (0.04) -0.81 (0.02) 0.66 (0.02) -0.27 (0.03)
1000 -8771 (137) 0.82 (0.03) -0.71 (0.02) 0.75 (0.01) -0.19 (0.02)
2000 -7953 (155) 0.83 (0.01) -0.34 (0.02) 0.68 (0.01) -0.08 (0.02)
@@
200 4.22 (0.31) 0.44 (0.06) 0.13 (0.01) 0.00 (0.00) -0.85 (0.09)
500 5.96 (0.56) -0.01 (0.06) 0.14 (0.01) 0.00 (0.00) -0.14 (0.01)
1000 8.05 (0.72) 0.16 (0.06) 0.26 (0.02) 0.00 (0.00) -0.17 (0.02)
2000 19.4 (1.87) 0.18 (0.05) 0.61 (0.02) 0.00 (0.00) -0.32 (0.02)